# VertiBench: Advancing Feature Distribution Diversity in Vertical Federated Learning Benchmarks

### v1

@@ 1 Introduction
The increasing demand for ample, high-quality data for training advanced machine learning models is evident, particularly in the context of large language models Brown et al. (2020). However, real data, often sensitive and distributed across multiple parties, presents a challenge, especially in the face of strict privacy regulations like the GDPR Voigt and Von dem Bussche (2017). As a result, federated learning Konečnỳ et al. (2016) has been highlighted as a promising approach to train machine learning models on distributed data while ensuring privacy. In this study, we consider a broad definition of federated learning Li et al. (2021a), encompassing all privacy-preserving collaborative learning paradigms, including assisted learning Diao et al. (2022a); Xian et al. (2020a) and split learning Vepakomma et al. (2018); Chen and Zhang (2022). Given the emerging variety of federated learning approaches Zhu et al. (2021); Li et al. (2021a); Yang et al. (2023), the importance of comprehensive benchmarks for evaluating these new algorithms is underscored.
The landscape of federated learning benchmarks, featuring contributions such as FedScale Lai et al. (2022), MNIST Lee and Shin (2020), FedEval Chai et al. (2020), and NIID-Bench Li et al. (2022a), predominantly caters to horizontal federated learning (HFL), wherein each party possesses a subset of instances. In comparison, vertical federated learning (VFL) - where each party holds a subset of features - is notably under-addressed.
The development of real-world VFL benchmarks faces two main hurdles. First, privacy concerns inherent to federated learning inhibit the public sharing of distributed data. Second, the current limited pool of actual VFL datasets, such as those in the OARF benchmark Hu et al. (2022), NUS-WIDE Chua et al. (July 8-10, 2009), and Vehicle Duarte and Hu (2004), may not sufficiently represent the broad range of possible VFL scenarios (Figure 1(c); Section 5). This scarcity and underrepresentation highlight the urgent need for synthetic VFL benchmarks that can facilitate a comprehensive evaluation of VFL algorithms across diverse scenarios.
Existing efforts to construct synthetic VFL benchmarks have struggled to represent the diversity of real-world scenarios. Benchmarks such as OARF Hu et al. (2022), FedML He et al. (2020), and LEAF Caldas et al. (2018) fabricate vertically partitioned data by randomly assigning an equal number of features to synthetic parties. Some studies Hu et al. (2019); Diao et al. (2022a); Castiglia et al. (2022a) resort to a simplistic approach, manually dividing features without offering a substantial rationale for their choice. Furthermore, these existing benchmarks Hu et al. (2022); He et al. (2020); Caldas et al. (2018) do not offer a meaningful comparison of cutting-edge VFL algorithms. Hence, it becomes crucial to critically examine the key factors influencing VFL algorithm performance, and thoughtfully design synthetic VFL benchmarks that reflect these considerations.
The task of creating a systematic synthetic VFL benchmark hinges on identifying key factors influencing VFL algorithm performance. Current synthetic benchmarks for non-i.i.d. HFL like NIID-Bench Li et al. (2022a) are inapplicable to VFL their assumptions about feature space and instance equality. In particular, HFL benchmarks assume that all parties operate within the same feature space, a presumption that doesn’t conform to VFL’s distributed feature scenario. Moreover, the instance equality presupposed by NIID-Bench during allocation doesn’t apply when dealing with features of varied importance, underscoring the unique challenges in the analysis of synthetic VFL benchmark.
Given these limitations, our study conducts a systematic analysis to identify feature importance and correlation as two crucial determinants of VFL performance. Accordingly, we propose VertiBench, a comprehensive VFL benchmark which introduces novel feature-splitting methods for synthetic dataset generation and a new real-world VFL dataset. Our key contributions include: (1) We develop new feature-splitting methods that generate synthetic datasets based on feature importance and correlation, covering a diverse range of VFL scenarios. (2) We introduce a real-world VFL dataset, filling a noted gap in image-image VFL scenarios. (3) We devise methods to quantify the importance and correlation of real-world VFL datasets, allowing them to align with our synthetic datasets. (4) We conduct rigorous benchmarks of advanced VFL algorithms across diverse scenarios, thereby offering valuable insights for future research. For example, we demonstrate the scalability of certain VFL algorithms, challenging prior assumptions about VFL scaling difficulties Hu et al. (2022), and emphasize the importance of communication efficiency in VFL, especially for imbalanced distributed datasets.
2 Related Work
Vertical federated learning datasets. The scarcity and limited range of real-world VFL datasets Chua et al. (July 8-10, 2009); Hu et al. (2022); Duarte and Hu (2004) in benchmarks and studies He et al. (2020); Caldas et al. (2018) underscore the need for synthetic VFL datasets capable of depicting a broader spectrum of scenarios. Given VFL’s focus on data privacy, obtaining such real datasets is challenging. Synthetic benchmarks He et al. (2020); Caldas et al. (2018) and VFL study datasets Diao et al. (2022a); Vepakomma et al. (2018); Cheng et al. (2021a) commonly rely on unexplained random or manual feature splitting, which often represents scenarios of balanced feature importance and high inter-party correlation (Figure 11(c)). This situation indicates a pressing demand for systematic methods to generate synthetic VFL datasets that accurately reflect a diverse set of scenarios, fostering comprehensive evaluation of VFL algorithms.
Feature importance. The Shapley value Cohen et al. (2005); Kumar et al. (2020); Ma and Tourani (2020), used to assess party contributions in federated learning Wang et al. (2019); Wang (2019); Han et al. (2021), has significant computational costs, rendering it unsuitable for guiding feature splitting. Certain methodologies Li et al. (2022a, 2023) utilize a Dirichlet distribution for global dataset random split, creating imbalanced federated learning datasets. However, they do not consider the partitioning of features of varying importance.
Feature correlation. The task of efficiently gauging correlation among two groups of features is challenging despite well-studied individual feature correlation Myers and Sirois (2004); De Winter et al. (2016). The Shapley-Taylor index, proposed for evaluating correlation between feature sets Sundararajan et al. (2020), is computationally intensive (NP-hard), and unsuitable for high-dimensional datasets. The determinant of the correlation matrix Wang and Zheng (2014) efficiently estimates inter-party correlation but is over-sensitive to linearly correlated features, impeding its use in feature partitioning. A more refined metric - the multi-way correlation coefficient (mcor) Taylor (2020), addresses this, but like the determinant, it struggles with unequal feature numbers across parties, a typical VFL scenario, due to the assumption of a square correlation matrix.
3 VFL Algorithms
This section critically reviews current VFL algorithms, with a focus on accuracy, efficiency, and communication size. VertiBench concentrates on standard supervised learning tasks such as classification and regression within synchronized parties, summarized in Table 1. Notably, this benchmark excludes studies exploring different VFL aspects such as privacy Jin et al. (2021), fairness Qi et al. (2022), data pricing Jiang et al. (2022), asynchronization Zhang et al. (2021a); Hu et al. (2019); Zhang et al. (2021b), latency Fu et al. (2022a), and other tasks like unsupervised learning Chang et al. (2020), matrix factorization Li et al. (2021b), multi-task learning Chen and Zhang (2022), and coreset construction Huang et al. (2022). While most VFL algorithms presume accurate inter-party data linking, we adopt this approach in VertiBench, despite recent contrary findings Wu et al. (2022); Nock et al. (2021) that this assumption may not be true. We refer to parties with and without labels as primary and secondary parties respectively.
Table 1: Summary of existing VFL algorithms
Category Model1 Algorithm Contribution2 Code3 Data4 Split5
Ensemble-based Any AL Xian et al. (2020a) Accuracy Xian et al. (2020b) Syn Manual
GAL Diao et al. (2022a) Accuracy Diao et al. (2022b) Syn Manual
Split-based NN SplitNN Vepakomma et al. (2018) Accuracy Wu (2022) Syn N/A
C-VFL Castiglia et al. (2022a) Communication Castiglia et al. (2022b) Syn Manual
BlindFL Fu et al. (2022b) Efficiency N/A Syn Manual
GBDT SecureBoost Cheng et al. (2021a) Accuracy Cheng et al. (2021b) Syn Manual
Pivot Wu et al. (2020a) Accuracy Wu et al. (2020b) Syn Manual
FedTree Li et al. (2023) Accuracy, Efficiency Li et al. (2022b) Syn Random
VF2Boost Fu et al. (2021) Efficiency N/A Syn Manual
  • 1 Abbreviations: NN - neural network; GBDT - gradient boosting decision trees; Any - model-agnostic.
  • 2 Privacy is not evaluated in VertiBench, and therefore omitted, despite being a contribution of some studies.
  • 3 Code availability: N/A - code not publicly accessible due to commercial issues.
  • 4 Dataset in experiments: Syn - synthetic datasets partitioned from global datasets.
  • 5 Datasets used in the experiments: Manual - features manually split without specific reasons; Random - features randomly split without explanation; N/A - no VFL experiments conducted.
The existing methods can be bifurcated into two categories: ensemble-based and split-based. The distinguishing factor lies in the independent prediction capability of each party. Ensemble-based methods involve parties each maintaining a full model for local feature prediction, with collaborative ensemble methods during training, while split-based methods require each party to hold a partial model forming different inference stages of the full model. Consequently, split-based partial models cannot perform independent inference. For split-based models, our focus is on advanced models such as neural networks (NNs) and gradient boosting decision trees (GBDTs) Chen and Guestrin (2016), though VertiBench can accommodate various models Hardy et al. (2017); Gu et al. (2020). Split-NN-based models are trained by transferring representations and gradients, while split-GBDT-models are trained by transferring gradients and histograms. A more detailed comparison of ensemble-based and split-based algorithms is provided in Appendix B.
4 Synthetic VFL Datasets
@@ 4.1 Factors that affect VFL performance
Suppose there are $K$ parties. Denote the data on party $P_{k}$ as a random vector $\mathbf{X}_{k}$ ($1\leq k\leq K$). Denote the label as a random variable $y$. A supervised learning algorithm maximizes the likelihood function where hypothesis $h$ represents models and parameters, i.e., $L(y|\mathbf{X}_{K},...,\mathbf{X}_{1};h)$.
These supervised learning algorithms estimate the following probability mass function. The proof of Proposition 1 is provided in Appendix A.
@@
$\log\mathcal{P}(y|\mathbf{X}_{K},...,\mathbf{X}_{1})=\sum_{i=1}^{K}\log{\frac{\mathcal{P}(y|\mathbf{X}_{k},...,\mathbf{X}_{1})}{\mathcal{P}(y|\mathbf{X}_{k-1},...,\mathbf{X}_{1})}}+\log\mathcal{P}(y)$ (1)
@@
In VFL, $\mathcal{P}(y)$ is the same for all the parties. The skewness among $K$ parties is determined by $K$ ratios of distributions. Interestingly, this ratio quantifies the divergence between two marginal probability distributions of $y$ one inclusive of $\mathbf{X}_{k}$ and the other exclusive of $\mathbf{X}_{k}$. Essentially, the ratio estimates the impact on the global distribution when the features of a single party are excluded. This can be interpreted as the importance of a given party.
It is important to note that Proposition 1 is applicable regardless of the order of $\mathbf{X}_{1},\dots,\mathbf{X}_{k}$. For a more precise evaluation of each party’s importance, especially considering the independence among features, Shapley value has proven to be a useful measure. It has been employed to estimate the importance of each party in vertical federated learning scenarios Wang et al. (2019); Han et al. (2021).
In another aspect, the ratio $\frac{\mathcal{P}(y|\mathbf{X}_{k},...,\mathbf{X}_{1})}{\mathcal{P}(y|\mathbf{X}_{k-1},...,\mathbf{X}_{1})}$ is determined by the correlation between $\mathbf{X}_{k}$ and $\mathbf{X}_{1},\dots,\mathbf{X}_{k-1}$. In other words, the global distribution is affected by the feature correlation between different parties.
In summary, we highlight feature importance and correlation as two crucial factors that could potentially influence the performance of VFL algorithms. We treat importance and correlation as independent variables affecting $\frac{\mathcal{P}(y|\mathbf{X}_{k},...,\mathbf{X}_{1})}{\mathcal{P}(y|\mathbf{X}_{k-1},...,\mathbf{X}_{1})}$ in our analysis, despite a potential innate correlation between the two. The subsequent sections will introduce our approach to generating synthetic datasets based by these two factors.
4.2 Feature Importance
In light of the computational expense incurred by the Shapley value method, an alternative and more efficient strategy is necessary to perform feature splits based on importance. With all parties exhibiting symmetry in the context of $\mathbf{X}$, varying the importance among parties essentially translates to varying the variance of the importance among them. Assuming each party $P_{i}$ possesses an importance factor $\alpha_{i}>0$, we propose the implementation of the Dirichlet distribution parameterized by $\{\alpha_{i}\}_{i=1}^{K}$ for feature splitting. This approach ensures two beneficial properties post-split: (1) a larger $\alpha_{i}$ guarantees a higher expected importance for $P_{i}$, and (2) a smaller $\|\{\alpha_{i}\}_{i=1}^{K}\|_{2}$ assures a greater variance in the importance among parties.
More specifically, we propose a feature splitting method based on feature importance. After initializing local datasets for each party, a series of probabilities $p_{1},\dots,p_{K}$ s.t. $\sum_{i=1}^{K}p_{i}=1$ is sampled from a Dirichlet distribution $\text{Dir}(\alpha_{1},\dots,\alpha_{K})$. Each feature is randomly allocated to a party $P_{k}$, selected based on the probabilities $p_{k}$. To accommodate algorithms that fail when faced with empty features, we can ensure each party is initially provided with a random feature before the algorithm is set in motion. Detailed formalization of this algorithm can be found in Appendix C.

### v2

@@ 1 Introduction
Federated learning Konečnỳ et al. (2016) is acknowledged for enabling model training on distributed data with enhanced privacy. In this study, we delve into the less explored vertical federated learning (VFL), where each party has a feature subset, aligning with a general definition of federated learning Li et al. (2021a) that includes privacy-preserving collaborative learning like assisted learning Diao et al. (2022) and split learning Vepakomma et al. (2018). The VFL application, depicted in Figure 0(a), involves an initial development phase using synthetic or real-world benchmarks, followed by deployment in actual federated environments upon validation.
Evaluating VFL algorithms is challenging due to the inherent confidentiality of VFL data Liu et al. (2022). The scope of party imbalance and correlation in existing real VFL datasets, termed the real scope, is limited. Datasets in the OARF benchmark Hu et al. (2022), FedAds Wei et al. (2023), NUS-WIDE Chua et al. (2009), and Vehicle Duarte and Hu (2004), predominantly represent scenarios where parties are balanced and exhibit weak correlations, as depicted in Figure 0(b).
To address the constraints inherent in the real scope, many VFL benchmarks Hu et al. (2022); He et al. (2020); Caldas et al. (2018) utilize synthetic datasets. This evaluation scope, termed uniform scope, represent the imbalance-correlation scope under an equal distribution of features among parties, either randomly or manually. The uniform scope, though commonly adopted in VFL experiments Diao et al. (2022); Castiglia et al. (2022), confines the evaluation to scenarios featuring balanced, strongly correlated parties according to Figure 0(b). Another critical limitation is the misalignment between the uniform scope and real scope, underscoring the imperative for a diverse and realistic VFL benchmark.
(a) Pipeline of VFL
(b) Estimated scope of VFL datasets
Figure 1: Overview of existing VFL piplines and datasets and the estimated scope of VFL datasets
Constructing a systematic synthetic VFL benchmark necessitates pinpointing the key factors affecting VFL algorithm performance. Existing synthetic benchmarks for non-i.i.d. horizontal federated learning (HFL), such as NIID-Bench Li et al. (2022a), fall short for VFL due to inherent assumptions about feature space and instance significance. Specifically, while HFL benchmarks typically assume independent and uniformly significant instances, this does not hold in VFL where features exhibit intrinsic correlations and differing importances. Furthermore, HFL benchmarks posit that all parties share the same feature space, a premise misaligned with VFL’s distributed feature paradigm. This delineates the unique analytical challenges inherent to synthetic VFL benchmarks.
Given these limitations, our statistical analysis of supervised VFL tasks identifies party importance and correlation as two crucial factors influencing target probability distributions in synthetic VFL datasets derived from the same global dataset. Accordingly, we propose VertiBench, a comprehensive VFL benchmark featuring novel feature-splitting methods for synthetic dataset generation. VertiBench offers three primary benefits: (1) it generally encompasses the uniform scope; (2) it effectively emulates the real scope, as evidenced by comparable performance on VertiBench-synthetic datasets; and (3) it introduces the capability to evaluate other scenarios that have not been explored in the previous studies, e.g. imbalanced feature split, broadening the scope of VFL evaluation.
Our primary contributions include: (1) Synthetic dataset generation methods with varied party importance and correlation, capturing a broad scope of VFL scenarios. (2) Novel real-world image-to-image VFL dataset Satellite. (3) Techniques to evaluate the party importance and correlation of real-world VFL datasets, enabling feature split comparison with synthetic VFL datasets. (4) Comprehensive benchmarks of mainstream cutting-edge VFL algorithms, providing key insights. For example, we demonstrate the scalability of VFL algorithms, challenging prior assumptions about VFL scaling difficulties Hu et al. (2022), and emphasize the challenges of communication efficiency in VFL datasets across varying imbalance levels. The VertiBench source code is accessible on GitHub [^1] https://github.com/Xtra-Computing/VertiBench and can be easily installed from PyPi [^2] https://pypi.org/project/vertibench/.
2 Evaluate VFL Datasets
In this section, our objective is to investigate the primary factors influencing VFL performance when generating synthetic VFL datasets from a fixed global dataset. Additionally, we explore methods to efficiently estimate these factors, guiding the subsequent feature split.
@@ 2.1 Factors that affect VFL performance
Suppose there are $K$ parties. Denote the data on party $P_{k}$ as a random vector $\mathbf{X}_{k}$ ($1\leq k\leq K$). Denote the label as a random variable $y$. A supervised learning algorithm maximizes the likelihood function where hypothesis $h$ represents models and parameters, i.e., $L(y|\mathbf{X}_{K},...,\mathbf{X}_{1};h)$. These supervised learning algorithms estimate the probability mass function in Eq. 1. The proof of Proposition 1 is provided in Appendix A.
@@
$\log\mathbb{P}(y|\mathbf{X}_{K},...,\mathbf{X}_{1})=\sum_{k=1}^{K}\log{\frac{\mathbb{P}(y|\mathbf{X}_{k},...,\mathbf{X}_{1})}{\mathbb{P}(y|\mathbf{X}_{k-1},...,\mathbf{X}_{1})}}+\log\mathbb{P}(y)$ (1)
@@
In VFL, $\mathbb{P}(y)$ is the same for all the parties. The skewness among $K$ parties is determined by $K$ ratios of distributions. Interestingly, this ratio quantifies the divergence between two marginal probability distributions of $y$ one inclusive of $\mathbf{X}_{k}$ and the other exclusive of $\mathbf{X}_{k}$. Essentially, the ratio estimates the impact on the global distribution when the features of a single party are excluded. This can be interpreted as the importance of a given party. Proposition 1 applies regardless of the order of $\mathbf{X}_{1},\dots,\mathbf{X}_{k}$. Shapley value, emphasizing feature independence, aids in precisely evaluating party importance in vertical federated learning, as demonstrated in Wang et al. (2019); Han et al. (2021).
In another aspect, the ratio $\frac{\mathbb{P}(y|\mathbf{X}_{k},...,\mathbf{X}_{1})}{\mathbb{P}(y|\mathbf{X}_{k-1},...,\mathbf{X}_{1})}$ is determined by the correlation between $\mathbf{X}_{k}$ and $\mathbf{X}_{1},\dots,\mathbf{X}_{k-1}$. In cases where the independence assumption underlying the Shapley value is invalidated, assessing each party’s impact on the global distribution becomes more accurate when based on feature correlation.
We identify feature importance and correlation as pivotal factors influencing VFL algorithm performance. For datasets with nearly independent features, the low inter-party correlation makes correlation-based splits less meaningful, suggesting the superiority of importance-based feature splits. Conversely, in datasets with highly correlated features, assessing individual feature importance becomes impractical, making correlation-based splits more suitable due to varying inter-party correlations.
Importance and correlation are treated as orthogonal evaluation factors applicable in distinct scenarios. While there may be an intrinsic link between them, our experiments indicate that focusing on one factor at a time yields explainable results reflective of real-world performance. As discussed in Appendix H, the interplay between importance and correlation can be complex. A joint optimization for both factors might be computationally intensive and less explainable, while providing limited additional insights. The subsequent sections will introduce our approach to evaluate these two factors and generating synthetic datasets based on each factor accordingly.
2.2 Evaluate Party Importance
To assess the importance for each party, we sum the importance of its features. While numerous methods to evaluate feature importance can be adopted in VertiBench, this study primarily focuses on two approaches: 1) Shapley Value: Feature importance is determined using Shapley values, efficiently estimated by evaluating the performance of a trained XGBoost Chen and Guestrin (2016) on random subsets. 2) Shapley-CMI Han et al. (2021): This approach, which does not rely on specific models, estimates the importance of each feature based on the Shapley-CMI applied to the global dataset. Both methods yield consistent and reasonable estimates of party importance.
2.3 Evaluate Party Correlation
The task of efficiently evaluating correlation among two groups of features is challenging despite well-studied individual feature correlation Myers and Sirois (2004); De Winter et al. (2016). The Shapley-Taylor index, proposed for evaluating correlation between feature sets Sundararajan et al. (2020), is computationally intensive (NP-hard), and unsuitable for high-dimensional datasets. The determinant of the correlation matrix Wang and Zheng (2014) efficiently estimates inter-party correlation but is over-sensitive to linearly correlated features, impeding its use in feature partitioning. A more refined metric - the multi-way correlation coefficient (mcor) Taylor (2020), addresses this, but like the determinant, it struggles with unequal feature numbers across parties, a typical VFL scenario, due to the assumption of a square correlation matrix.
Given the limitations of existing metrics Taylor (2020); Wang and Zheng (2014), we propose a novel metric to examine the correlation when the parties involved possess unequal numbers of features. Our approach hinges on the use of the standard variance of the singular values of the correlation matrix. This serves as an efficient measure of the overall correlation between two parties. Since the feature-wise correlation is an orthogonal research area, we selected Spearman rank correlation Zar (2005) due to its capability to handle non-linear correlation.
To elaborate further, we denote the column-wise correlation matrix between two matrices, $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$, as $\text{cor}(\mathbf{X}_{i},\mathbf{X}_{j})$. As a result, we formally define the correlation between two entities, $\mathbf{X}_{i}\in\mathbb{R}^{n\times m_{i}}$ and $\mathbf{X}_{j}\in\mathbb{R}^{n\times m_{j}}$, in terms of their respective parties as Eq. 2.
$\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{j}):=\frac{1}{\sqrt{d}}\sqrt{\frac{1}{{d-1}}\sum_{t=1}^{d}{\left(\sigma_{t}(\text{cor}(\mathbf{X}_{i},\mathbf{X}_{j}))-\overline{\sigma}\right)^{2}}},\quad d=\min(m_{i},m_{j})$ (2)
In this equation, $\sigma_{i}(\cdot)$ means the $i$-th singular value of a matrix, while $\overline{\sigma}$ stands for their mean value. Proposition 2 states that Pcor is equivalent to mcor for inner-party correlation (see Appendix A for proof). Experiments detailed in Appendix D.1 reveal that Pcor exhibits trends analogous to mcor Taylor (2020) when assessing inter-party correlation between equal number of features.
Proposition 2.
[…]
The singular values of a correlation matrix, Pcor, represent the magnitudes of its ellipsoid’s semi-axes, indicating the degree of dependence among features. The standard deviation of these singular values reflects the distribution of dependence across different axes. A notably large singular value in a specific axis (Figure 1(c)) suggests a high concentration of dependence. For instance, if there’s only one nonzero singular value, it implies that all features are perfectly correlated with a single feature. Conversely, if the singular values are uniformly distributed such as Figure 1(a) (indicated by a small standard deviation), it denotes less concentrated feature correlations. Therefore, the standard deviation of singular values serves as a measure of the dataset’s proximity to perfect correlation.
Proposition 3 states that Pcor, like mcor, spans a range from 0 to 1, even when assessing inter-party correlation. A Pcor value of 1 signifies perfect correlation between $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, while a value of 0 indicates their independence.
Proposition 3.
[…]
It is important to note that the absolute value of Pcor alone does not fully capture inter-party correlation. For instance, when $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$ are two parties both containing the same set of independent features, $\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{j})$ yields a value of 0, the same as the Pcor between two independent parties. Despite the same Pcor value, these scenarios intuitively differ in their levels of inter-party correlation. This discrepancy arises from overlooking the inner-party correlation of $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$. Typically, parties with highly correlated features tend to exhibit higher Pcor values with other parties.
To accurately measure the correlation between $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$, we evaluate how the shift towards perfect correlation varies when $\mathbf{X}_{i}$ is replaced by $\mathbf{X}_{j}$. This is captured by the relative change in Pcor, denoted as $\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{j})-\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{i})$. In the perspective of variance analysis Kruskal and Wallis (1952), this difference quantifies the degree to which the standard deviation $\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{j})$ is explained by inter-party factors, controlling the contribution of inner-party correlations. The overall inter-party correlation, denoted as Icor, is described as the mean party-wise correlation across all distinct party pairs. Formally,
$\text{Icor}(\mathbf{X}_{1},\dots,\mathbf{X}_{K}):=\frac{1}{K(K-1)}\sum_{i=1}^{K}\sum_{j=1,j\neq i}^{K}{\left(\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{j})-\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{i})\right)}.$ (3)
(a) $x,y,z\sim U(0,1)$
(b) $x,y\sim U(0,1),z=-x^{2}-y^{2}$
(c) $x\sim U(0,1),y=2x,z=x+1$
Figure 2: Examples of Pcor values on different levels of correlation. $U$ means uniform distribution. Arrow direction indicates right singular vector orientation, arrow scale represents singular values.
Icor exhibits notable properties both theoretically and empirically. Theoretically, as demonstrated in Theorem 1 (see Appendix A for proof), optimizing Icor yields ideal feature splits in optimal scenarios. Specifically, in datasets comprising two independent but internally perfectly correlated feature sets, Icor reaches its minimum when each party exclusively possesses one feature set and attains its maximum when each party equally shares half of the features from both sets. Empirically, we evaluate the link between inter-party correlation and Icor in complex, real-world datasets (Appendix D). These empirical observations align with theoretical insights, confirming Icor’s capability in analyzing intricate data correlations.